proof that (x1 × d ln f1/d x1) + (x2 × d ln f2/d x1)=0 in chemical engineering thermodynamics
1
Expert's answer
2020-03-25T12:46:55-0400

"\phi=\sum\ x_i ln\ \hat{\phi_i}" (As per standard notations).

Here "\phi" is a molar property with "ln\ \hat{\phi_i}" be the property of a component in solution.

By using Gibbs-Duhem equation,

At constant "p,"

"x_1\ (d\ ln\hat{\phi_1})+x_2\ (d\ ln\hat{\phi_2})=0"

And,"\hat{f_i} =x_i \hat{\phi_i}{p}\implies \hat{\phi_i}=\frac{\hat{f_i}}{x_i \ p}"

"\implies x_1\ (d\ ln\frac{\hat{f_1}}{p\ x_1})+x_2\ (d\ ln\frac{\hat{f_2}}{p \ x_2})=0"

"\implies x_1\ ( d(\ ln{\hat{f_1}})-d(ln{\ x_1 p}))" "+x_2(\ (d\ ln{\hat{f_2}})-d(ln \ { x_2p\ }))=0"

As pressure is constant,"d(ln \ { x_1p\ })=d(ln \ { x_2p\ })=0"

The equation reduces to "x_1(d\ ln{\hat{f_1}})+x_2(d\ ln{\hat{f_2}})=0"

By using Lewis-Randall Rule,"\hat{f_1}=x_1f_1" ,

We get, "x_1\times d\ ln{(\ x_1{f_1)}}+x_2\times d\ ln{{\ (x_2 f_2)}}=0"